OpenStudy (anonymous):

4x2 + 12x = 0 a) x = 0, -3 b) x = 0, 3 c) x = -3 d) x = 3

OpenStudy (anonymous):

a

OpenStudy (anonymous):

a) x = 0, -3

OpenStudy (anonymous):

4x(x+3) = 0 if 4x = 0, then x = 0/4 = 0 if x+3 = 0, then x= -3 so 0, -3